# Causality-Respecting Adaptive Refinement for PINNs: Enabling Precise Interface Evolution in Phase Field Modeling

### v1

@@ 3 Phase Field modeling employing Allen-Cahn equation
[…]
Here, $W$ represents the energy barrier and $\kappa$ denotes the scale factor of the interfacial energy density. The function $g(u)=u^{2}(1-u)^{2}$ is the double well function, a fundamental component in phase field modeling [27, 28]. To describe the phase evolution, we employ the functional derivative of the Helmholtz free energy. Consequently, the corresponding Allen-Cahn equation for this system is given by:
@@
where $L_{\sigma}$ denotes the relaxation coefficient, which is proportional to the rate of phase evolution [29]. It is important to note that Eq. (11) describes a static co-existence of two phases with a finite-thickness interface [30]. To extend the model to dynamic systems with evolving phases, we introduce an additional driving force term to the static PFM:
$\frac{\partial u}{\partial t}$ $=-L_{\sigma}\left[2Wu(1-u)(1-2u)-\kappa\nabla^{2}u\right]+L_{\eta}\left[\frac{dP(u)}{du}f(u,\bm{x},t)\right]$ (12)
@@
In this extended form, $L_{\eta}$ represents the scaling coefficient for the driving force term, and $f(u,\bm{x},t)$ is the driving force function depending on the order parameter $u$, spatial coordinates $\bm{x}$, and time $t$. For simplicity in this study, we assume $f(u,\bm{x},t)$ to be a constant $c$, representing a constant driving force. The term $\frac{dP(u)}{du}=u^{3}(10-15u+6u^{2})$ ensures that the driving force originates from the interface, as proposed by Wang et al. [31]. This formulation provides a comprehensive framework for modeling phase evolution in both static and dynamic contexts, capturing the interplay between interfacial energy, bulk free energy, and external driving forces.
@@ 4 Numerical results
In this section, we employ the proposed RBAR causality-based PINN in both static and dynamic cases with different initial morphology, i.e., plane interface and hump interface, respectively. The static and the dynamic cases are denoted with Eq. (11) and (12), respectively. The solutions obtained via our PINN framework are systematically benchmarked against finite element method (FEM) solution computed using COMSOL Multiphysics. Through comprehensive comparative analysis, we demonstrate the computational efficiency, numerical accuracy, and adaptive error correction capabilities of the RBAR causality-based PINN methodology.
The neural network architecture employed in this work comprises a feed forward fully connected network with a hyperbolic tangent activation function, including 5 hidden layers and 100 neurons in each hidden layer.
@@ 4.1 Quasi-static Allen-Cahn Equation
We conduct our numerical experiments in a three-dimensional spatio-temporal domain where $(t,x,y)\in[0,1]\times[0,1]\times[0,1]$. The initial discretization comprises 8000 uniformly distributed collocation points prior to refinement, as depicted in Figure 2(a). Our two-phase PFM simulation tracks the phase transition dynamics, with the full three-dimensional evolution shown in Figure 2(b). The static Allen-Cahn equation takes the form in Eq. (11). And we set the corresponding parameters: $L_{\sigma}=1$, $W=8\times 10^{-4}$, and $\kappa=1\times 10^{-4}$. The initial condition, illustrated in Figure 1(d), establishes a discrete plane interface at $x=0.5$ separating two phases: the red phase ($u=1$) and the blue phase ($u=0$). Following observations were made in the initial implementation of the framework which employs only RBAR (without causality training):
@@
    The mean loss using the Adam optimizer [32] converges to below $10^{-10}$ (see Figure 2(c)), yet significant discrepancies emerge between PINN predictions (see Figures 2(d-f)) and COMSOL FEM results (see Figures 2(g-i)).
@@
Figure 2: Static phase evolution results with planar interface configuration. (a) Initial computational domain discretized with 8,000 ($20\times 20\times 20$) uniformly distributed collocation points. (b) Three-dimensional phase morphology evolution simulated using the PINN framework. (c) History of the mean loss obtained using the Adam optimizer over 20,000 training iterations. (d-f) Phase morphology evolution predicted by PINN at t = 0 s, 0.5 s, and 1 s, respectively. (g-i) Corresponding phase morphology evolution computed using COMSOL FEM simulation at identical time points for direct comparison. Color scale represents phase parameter u, where red (u = 1) and blue (u = 0) indicate distinct phases.
Following a single iteration of the computational domain with the collocation points in Figure 2(a), an adaptive refinement was implemented employing the RBAR method, as shown in Figure 3(a). The refinement process automatically identified regions of highest residual loss, which predominantly occurred along the sharp phase interface. This adaptive point distribution proved crucial for several reasons:
@@
Figure 3: Static phase evolution results with planar interface using RBAR-enhanced PINN methodology. (a) Computational domain after RBAR refinement showing adaptive point distribution. (b) Three-dimensional visualization of phase morphology evolution computed using RBAR PINN. (c) Mean Adam loss evolution per collocation point over 100,000 total iterations, comprising 50,000 iterations pre-refinement and 50,000 post-RBAR implementation. (d-f) Phase morphology evolution captured by RBAR PINN at t = 0 s, 0.5 s, and 1 s, demonstrating interface dynamics.
@@ 4.2 Dynamic Allen-Cahn Equation
where $W=8\times 10^{-4}$, $\kappa=1\times 10^{-4}$, $L_{\eta}=0.01$, and $c=200$. The initial condition consists of a planar interface morphology. In this investigation of moving plane dynamics, we implement RBAR PINN without the causality training strategy. Impact of Refinement Strategies. Our observations are summarized as follows:
@@
RBAR Implementation: Our optimal strategy consisted of an initial training phase with 50,000 iterations, followed by RBAR-guided domain refinement (illustrated in Fig. 4) which is the secondary training phase that has an additional 50,000 training iterations.
@@ 4.3 Simulating complex Morphology with Dynamic Equation
Limitations of RBAR-Only Implementation: Initial attempts using only RBAR refinement revealed significant challenges in accurately capturing the hump evolution. While multiple refinement iterations improved certain aspects of the solution—including better interface smoothness and phase-field values more strictly confined to the $[0,1]$ range—the predicted morphology deviated substantially from the COMSOL reference solution. Specifically, the characteristic hump structure disappeared as the interface evolved toward a planar configuration, indicating fundamental limitations of the RBAR-only approach for this dynamic case.
Figure 6: Evolution of phase-field solution using RBAR-causality combined PINN methodology: comparison of results with no refinement, first, second, and third refinement iterations, showing progressive improvement in accuracy.
@@
To elucidate the complementary roles of RBAR and causality training, we conducted detailed analysis of the early-time evolution (see Figure 7). The causality-training PINN accurately captures the initial morphology dynamics, matching COMSOL results in both interface position and hump structure. RBAR then identifies and refines regions of highest residual loss—primarily the interface and hump regions—providing enhanced spatial resolution where needed most. This synergistic interaction enables accurate capture of complex morphological features, elimination of spatial overshooting as well as correct positioning of the evolving interface
The combined methodology successfully reproduces the COMSOL reference solution, demonstrating the effectiveness of integrating RBAR refinement with causality training for challenging phase-field evolution problems.
Figure 7: Comparative analysis of early-stage evolution (t = 0s, 0.1s, 0.2s) showing phase-field morphology obtained via COMSOL, causality training without refinement, and RBAR refinement patterns.
@@ 5 Conclusion
In this study, we introduce a novel methodology that combines RBAR (Residual-Based Adaptive Refinement) with causality-informed training in PINNs to solve spatio-temporal PDEs, with specific application to the phase field model (PFM) based on Allen-Cahn equations. Our analysis of the dynamic case with planar interfaces reveals that uniform refinement of the computational domain proves inefficient. Instead, the RBAR mechanism enables the PINN to overcome convergence to erroneous solutions by strategically allocating collocation points to capture interface evolution phenomena. However, we observed that RBAR alone proves insufficient for accurately resolving dynamic PFM cases with complex interfacial morphologies. The integration of causality-informed training with RBAR significantly enhances the framework’s capability to accurately reproduce temporal interface evolution, particularly in cases involving non-planar morphologies. During the iterative refinement and training process, we identify a characteristic overshoot and relocate phenomenon, which emerges from the synergistic interaction between RBAR and causality-informed training. This mechanism operates as follows: initially, RBAR concentrates collocation points in regions of high loss that have been partially resolved through causality training. Subsequently, the causality-informed training utilizes these strategically placed points to obtain accurate solutions at initial timesteps. These refined initial solutions then enable accurate prediction and tracking of interface evolution in subsequent timesteps.
While our current implementation focuses on dynamic Allen-Cahn equations, the proposed RBAR-causality framework is readily extensible to diverse spatio-temporal systems due to its generalized formulation. The RBAR methodology demonstrates simultaneous improvements in solution accuracy and computational efficiency. Furthermore, this combined RBAR causality-based PINN framework establishes a foundational methodology for future developments in PINN-based solutions of Phase Field Models.
@@ Acknowledgements
The authors (WW, TPW, and HHR) would like to acknowledge the support by the Hong Kong General Research Fund (GRF) under Grant Numbers 15213619 and 15210622, and by an industry collaboration project (HKPolyU Project ID: P0039303). SG is supported by U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research grant under Award Number DE-SC0024162 and the 2024 Johns Hopkins Discovery Award. The authors acknowledge the computational resources provided by the University Research Facility in Big Data Analytics (UBDA) at The Hong Kong Polytechnic University.
@@
The authors declare no competeting interest

### v2

@@ 3 Phase Field modeling employing Allen-Cahn equation
$\mathcal{F}=\int_{\Omega}fd\omega=\int_{\Omega}(Wg(u)+\frac{\kappa}{2}|\nabla u|^{2})d\omega.$ (12)
Here, $W$ represents the energy barrier and $\kappa$ denotes the scale factor of the interfacial energy density. The function $g(u)=u^{2}(1-u)^{2}$ is the double well function, a fundamental component in phase field modeling Lin and Ruan [2019], Moelans et al. [2008]. To describe the phase evolution, we employ the functional derivative of the Helmholtz free energy. Consequently, the corresponding Allen-Cahn equation for this system is given by:
@@
where $L_{\sigma}$ denotes the relaxation coefficient, which is proportional to the rate of phase evolution Fan and Chen [1997]. It is important to note that Eq. (13) describes a static co-existence of two phases with a finite-thickness interface Kim et al. [1999]. To extend the model to dynamic systems with evolving phases, we introduce an additional driving force term to the static PFM:
$\frac{\partial u}{\partial t}$ $=-L_{\sigma}\left[2Wu(1-u)(1-2u)-\kappa\nabla^{2}u\right]+L_{\eta}\left[\frac{dP(u)}{du}f(u,\bm{x},t)\right]\$(14)
@@
In this extended form, $L_{\eta}$ represents the scaling coefficient for the driving force term, and $f(u,\bm{x},t)$ is the driving force function depending on the order parameter $u$, spatial coordinates $\bm{x}$, and time $t$. For simplicity in this study, we assume $f(u,\bm{x},t)$ to be a constant $c$, representing a constant driving force. The term $\frac{dP(u)}{du}=u^{3}(10-15u+6u^{2})$ ensures that the driving force originates from the interface, as proposed by Wang et al. Wang et al. [1993]. This formulation provides a comprehensive framework for modeling phase evolution in both static and dynamic contexts, capturing the interplay between interfacial energy, bulk free energy, and external driving forces.
@@ 4 Numerical results
In this section, we employ the proposed RBAR causality-based PINN in both static and dynamic cases with different initial morphology, i.e., plane interface and hump interface, respectively. The static and the dynamic cases are denoted with Eq. (13) and (14), respectively. The solutions obtained via our PINN framework are systematically benchmarked against the FEM solution computed using COMSOL Multiphysics. Through comprehensive comparative analysis, we demonstrate the computational efficiency, numerical accuracy, and adaptive error correction capabilities of the RBAR causality-based PINN methodology in comparison with the conventional PINN schemes.
@@ 4.1 Quasi-static Allen-Cahn Equation
We conduct our numerical experiments in a three-dimensional spatio-temporal domain where $(t,x,y)\in[0,1]\times[0,1]\times[0,1]$. The initial discretization comprises 8000 uniformly distributed collocation points prior to refinement, as depicted in Figure 2(a). Our two-phase PFM simulation tracks the phase transition dynamics, with the full three-dimensional evolution shown in Figure 2(b). The static Allen-Cahn equation takes the form in Eq. (13). And we set the corresponding parameters: $L_{\sigma}=1\;m^{3}(J\cdot s)^{-1}$, $W=8\times 10^{-4}\;J\cdot m^{-3}$, and $\kappa=1\times 10^{-4}\;J\cdot m^{-1}$. The initial condition, illustrated in Figure 1(d), establishes a discrete plane interface at $x=0.5$ separating two phases: the red phase ($u=1$) and the blue phase ($u=0$). The following observations were made in the initial implementation of the framework, which employs only RBAR (without causality training):
@@
    The mean loss using the Adam optimizer Kingma [2014] converges to below $10^{-10}$ (see Figure 2(c)), yet significant discrepancies emerge between PINN predictions (see Figures 2(d-f)) and COMSOL FEM results (see Figures 2(g-i)).
@@
Figure 2: Static phase evolution with planar interface using conventional PINN. (a) Initial computational domain discretized with 8,000 ($20\times 20\times 20$) uniformly distributed collocation points. (b) Three-dimensional evolution of the phase morphology simulated using the PINN framework. (c) History of the mean training loss obtained with the Adam optimizer over 20,000 iterations. (d–f) Phase morphology evolution predicted by PINN without RBAR refinement at $t=0$, $0.5$, and 11 s, respectively. (g–i) Corresponding phase morphology evolution computed using COMSOL FEM at the same time points for direct comparison. The color scale represents the phase parameter $u$, with red ($u=1$) and blue ($u=0$) denoting distinct phases.
Following a single iteration of the computational domain with the collocation points in Figure 2(a), an adaptive refinement was implemented employing the RBAR method, as shown in Figure 3(a). The refinement process automatically identified regions of highest residual loss, which predominantly occurred along the sharp phase interface. While some of the refined points appear to be randomly distributed, resulting in an irregular and seemingly counterintuitive pattern, this behavior stems from the inherently non-linear optimization process of PINNs. Factors such as training dynamics, sensitivity to initialization, and convergence to local minima can introduce significant errors in unexpected regions of the domain, thereby leading to a degree of randomness in the refinement pattern. This adaptive point distribution proved crucial for several reasons:
@@
Figure 3: Static phase evolution with planar interface using the RBAR-PINN methodology. (a) Computational domain after RBAR refinement, showing the adaptive distribution of collocation points. (b) Three-dimensional evolution of the phase morphology predicted by RBAR-PINN. (c) Mean training loss per collocation point using the Adam optimizer over 100,000 iterations, including 50,000 pre-refinement and 50,000 post-RBAR iterations. (d–f) Phase morphology evolution captured by RBAR-PINN at $t=0$, $0.5$, and 11 s, illustrating the interface dynamics.
@@ 4.2 Dynamic Allen-Cahn Equation
where $W=8\times 10^{-4}\;J\cdot m^{-3}$, $\kappa=1\times 10^{-4}\;J\cdot m^{-1}$, $L_{\eta}=0.01\;s^{-1}$, and $c=200$. The initial condition consists of a planar interface morphology. In this investigation of moving plane dynamics, we implement RBAR PINN without the causality training strategy. Impact of Refinement Strategies. Our observations are summarized as follows:
@@
RBAR Implementation: Our optimal strategy consisted of an initial training phase with 50,000 iterations, followed by RBAR-guided domain refinement (illustrated in Fig. 4), which is the secondary training phase that has an additional 50,000 training iterations.
@@ 4.3 Simulating complex Morphology with Dynamic Equation
Limitations of RBAR-Only Implementation: Initial attempts using only RBAR revealed significant challenges in accurately capturing the hump evolution. While multiple refinement iterations improved certain aspects of the solution—including better interface smoothness and phase-field values more strictly confined to the $[0,1]$ range—the predicted morphology deviated substantially from the COMSOL reference solution. Specifically, the characteristic hump structure disappeared as the interface evolved toward a planar configuration, indicating fundamental limitations of the RBAR-only approach for this dynamic case.
Figure 6: Evolution of phase-field solution using RBAR-causality combined PINN methodology: comparison of results with no refinement, first, second, and third refinement iterations, showing progressive improvement in accuracy. The last row presents the absolute error between the three-times-refined PINN predictions and the reference solutions obtained from COMSOL.
@@
The last row of Figure 6 shows the absolute error between the three-times-refined RBAR-PINN predictions and the reference solutions obtained from COMSOL. The error is highly localized at the interface and remains negligible across the rest of the domain.
To elucidate the complementary roles of RBAR and causality training, we conducted detailed analysis of the early-time evolution (see Figure 7). It should be emphasized that the refinement process yields approximately 4,000 points per time step, which is ten times the original count of 400 points. The causality-training PINN accurately captures the initial morphology dynamics, matching COMSOL results in both interface position and hump structure. RBAR then identifies and refines regions of highest residual loss—primarily the interface and hump regions—providing enhanced spatial resolution where needed most. This synergistic interaction enables accurate capture of complex morphological features, elimination of spatial overshooting, as well as correct positioning of the evolving interface
The combined methodology successfully reproduces the COMSOL reference solution, demonstrating the effectiveness of integrating RBAR with causality training for challenging phase-field evolution problems. While the well-established finite element software COMSOL demonstrates significantly greater computational efficiency than the RBAR-causality combined PINNs, completing the simulation in approximately 20 seconds compared to 3 hours. It is important to note that physics-informed neural networks are a nascent, mesh-free class of methods for solving PDEs, and they face well-recognized challenges when modeling systems with sharp, evolving interfaces. The objective of the present work is therefore not to compete with mature finite element solvers in terms of raw computational efficiency, but to address a critical failure mode of conventional PINNs. The effectiveness and efficiency of the proposed RBAR–causality combined framework are demonstrated specifically through systematic comparison with traditional PINN formulations, establishing a meaningful advancement for PINN-based modeling of complex interface-evolving systems.
Figure 7: Comparative analysis of early-stage evolution (t = 0s, 0.1s, 0.2s) showing phase-field morphology obtained via COMSOL, causality training without refinement, and RBAR patterns.
@@ 5 Conclusion
In this study, we introduce a novel methodology that combines RBAR (Residual-Based Adaptive Refinement) with causality-informed training in PINNs to solve spatio-temporal PDEs, with a specific application to the phase field model (PFM) based on Allen-Cahn equations. Our analysis of the dynamic case with planar interfaces reveals that uniform refinement of the computational domain proves inefficient. Instead, the RBAR mechanism enables the PINN to overcome convergence to erroneous solutions by strategically allocating collocation points to capture interface evolution phenomena. However, we observed that RBAR alone proves insufficient for accurately resolving dynamic PFM cases with complex interfacial morphologies. The integration of causality-informed training with RBAR significantly enhances the framework’s capability to accurately reproduce temporal interface evolution, particularly in cases involving non-planar morphologies. During the iterative refinement and training process, we identify a characteristic overshoot and relocate phenomenon, which emerges from the synergistic interaction between RBAR and causality-informed training. This mechanism operates as follows: initially, RBAR concentrates collocation points in regions of high loss that have been partially resolved through causality training. Subsequently, the causality-informed training utilizes these strategically placed points to obtain accurate solutions at initial timesteps. These refined initial solutions then enable accurate prediction and tracking of interface evolution in subsequent timesteps.
Finally, it should be emphasized that while this study focuses on the benchmark Allen-Cahn equation with fixed parameters, the RBAR-PINN framework is fundamentally designed to be parameter-agnostic. By successfully resolving the numerical“stiffness” associated with sharp, moving interfaces, a known failure point for standard PINN formulations, this work establishes a robust computational foundation for dynamic phase-field modeling. The ability of the algorithm to adaptively redistribute collocation points based on residual ranking ensures that the method can accommodate varying material properties, such as mobility and surface energy, which dictate different interface widths and velocities. Future research will leverage this validated framework to investigate multi-component systems and a broader range of material scales, moving beyond proof-of-concept benchmarks toward complex industrial applications in materials science.
@@ Acknowledgements
The authors (WW, TPW, and HHR) would like to acknowledge the support by the Hong Kong General Research Fund (GRF) under Grant Numbers 15213619 and 15210622, by an industry collaboration project (HKPolyU Project ID: P0039303), and by a collaborative research fund from HKPolyU (Project ID: P0058676). SG is supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research grant under Award Number DE-SC0024162 and the 2024 Johns Hopkins Discovery Award. The authors acknowledge the computational resources provided by the University Research Facility in Big Data Analytics (UBDA) at The Hong Kong Polytechnic University.
@@
The authors declare no competing interest